OpenStudy (anonymous):

Marginal utility of x/Marginal utility of y= (2x^(-5/6) . y^(1/6))/(12x^(1/6) . (1/6)y^(-5/6))= Price x/ price y please help me solve for x

OpenStudy (anonymous):

i have \[(12x^\frac{ 1 }{ 6 }.\frac{ 1 }{ 6 } .y^\frac{ -5 }{ 6 } ) . \frac{ price x }{ price y } = 2x^\frac{ -5 }{ 6 } . y ^\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

then \[((12x^\frac{ 1 }{ 6 }.\frac{ 1 }{ 6 } .y^\frac{ -5 }{ 6 } ) . \frac{ price x }{ price y })/y \frac{ 1 }{ 6 } = 2x^\frac{ -5 }{ 6 }\]

OpenStudy (anonymous):

but now i m stuck and dont know how to go further